03 March 2013

A successful first week

The first week of the 2013 campaign has come to a successful close. The weather has been fantastic for solar measurements, thanks to an unusually high pressure system above Eureka. Since today was Sunday, we left the Eureka Weather Station after brunch and arrived at PEARL shortly before noon. The Sun was bright, the sky was extraordinarily clear, and wind was minimal. The temperature at the Ridge Lab was wonderfully warm, at -28°C. Eureka remains at -48°C. The Ridge Lab team stopped briefly along the drive up to watch Arctic Hares near the side of the road.

In the PEARL Ridge Lab's IR lab, Joseph and Dan took 46 MIR measurements with the Bruker; Debora took 25 measurements with PARIS.

In the UV-Vis Lab, Xiaoyi has kept the UT-GBS working with a diffuser on its telescope. SAOZ continues to operate normally. Xiaoyi installed a new wedged CCD mount for the PEARL-GBS, and performed standard tests to check the resolution, stray light, and polarization. At the end of the day, he moved the PEARL-GBS telescope back into the mezzanine and confirmed that the spectra continue to look good. Xiaoyi left the PEARL-GBS taking dark current measurements overnight.

Volodya took Brewer #069 inside for servicing today. Brewer #192 stopped working again. It had been working normally, but may require another trip inside for inspection. MAESTRO and SPS are working well. Their data is being copied to CANDAC servers daily.

Down at 0PAL, Zen monitored the E-AERI status, and did ëspring cleaning' inside the 0PAL container. There is a water leakage problem with the thru-wall (between the two sections of the E-AERI). Unfortunately, he will have to wait until it rains in the spring to find exactly where the water is leaking in from.

Chris remotely performed CRL measurements 24 hours a day. There was a minor interruption of 10 minutes today, but the lidar was brought back online with no issues. Emily tested the laser burn quality; Laser #1 continues to perform well. Emily also disassembled the spare pump chamber with the intention of replacing the o-rings around the laser rod and replacing the flashlamps. Although the rod is in good shape, she unfortunately found evidence of corrosion on the internal surfaces of the pump chamber. The CRL group is discussing this.

Today's ACE ozonesonde flight occurred at 18:34 (local time) and reached an altitude of 21,637 m (37.5 hPa).

After a long week of hard work and great success, the team relaxed with a movie at the Eureka Weather Station tonight.

Dan

Back to Daily Reports
Back to Home